Maura ran 3/5 of the course and walked the rest. How much of the course did she walk?

Answer by Schaman_Dempster(26)

Let the distance walked be x.
The total course would be 1 (1/1)
So, distance ran + distance walked = total course length
3/5 + x = 1
x = 1 - 3/5
x = 5/5 - 3/5
x = (5-3)/5
x = 2/5
She walked 2/5 of the course.